# Angular-momentum pairs in spherical systems: applications to the Galactic centre

### v1

@@ 5 Applications to the Galactic Centre
We run a set of isolated three-ring problems in N-Ring with initial conditions adjusted to mimic the Galactic centre. We initialise all pairs with relative inclinations of $\theta_{b}=10^{\circ}$ and inclination with respect to the perturber $\theta_{\bullet}\simeq 65^{\circ}$. We choose this inclination because, in the case of disruption, the maximum relative angle $\theta_{1}+\theta_{2}$ exceeds $100^{\circ}$, pushing the corresponding pairs into retrograde orbits relative to the inner objects. For all pairs, we set eccentricities $e_{s}=0.3\pm 0.03$, and we explore a range of semi-major axes $0.04\leq a\leq 0.5~{}\text{pc}$. This parameter set mimics that of the young stars in the central parsec of the Galaxy (von Fellenberg et al., 2022). For the perturber, we set the eccentricity to $e_{\bullet}=0.3$ but vary its semi-major axis. The discrete values chosen are: $a_{\bullet}=0.15$ pc, the projected distance of the IRS13 association, potentially hosting an IMBH (Tsuboi et al., 2017); $a_{\bullet}=0.5$ pc, the possible 3D distance of IRS13 (Tsuboi et al., 2020); and $a_{\bullet}=1.5$ pc, an estimated distance to the circumnuclear disc (e.g. Oka et al. 2011). Additionally, we vary the mass of the perturber from $2\times 10^{3}$ to $10^{5}~{}\,{\rm M}_{\odot}$, placing lighter perturbers closer to the SMBH to operate within the allowed range of masses for the putative IMBHs (Lommen & Backer, 2001; Gualandris & Merritt, 2009; Gualandris et al., 2010; Naoz et al., 2020). For each set of initial conditions, we perform $10^{4}$ three-ring simulations.
@@
To interpret the results, we stack all pairs in a single snapshot at $t=10~{}\text{Myr}$ and measure their relative inclinations $\theta_{b}$, as well as the relative inclination $\theta_{\mathrm{rel}}$ of each pair’s total angular momentum $\bm{L}_{i}+\bm{L}_{j}$ with respect to the net angular momentum of the innermost pairs (semi-major axes $0.04\leq a\leq 0.05$ pc). An important parameter is the mass of the fragment relative to the IMBH mass; we thus choose two different clump masses: $m_{i}+m_{j}=100\,{\rm M}_{\odot}$ and $m_{i}+m_{j}=200\,{\rm M}_{\odot}$, with $m_{i}=m_{j}$ in both cases.
@@
[…]
@@
In summary, our model favours an IMBH mass of about $5\times 10^{3}\,{\rm M}_{\odot}$. However, this model treats the disc-fragment as two rings, neglecting its interaction with the rest of the disc, and also does not include torques due to fluctuating anisotropy in the nuclear star cluster, which could be significant on the relevant time-scales (e.g. Panamarev & Kocsis 2022), nor does it account for the inward migration of the IMBH induced by Chandrasekhar dynamical friction (Chandrasekhar, 1943), or an orbital alignment due to resonant dynamical friction (Szölgyén et al., 2021; Ginat et al., 2023). Including these additional effects may reveal that lower-mass IMBHs can also effectively reorient the stellar disc into retrograde orbits within $\sim 10$ Myr. Although our simplified model provides useful order-of-magnitude estimates, it can neither rule out the existence of IMBHs, nor prove them – merely point to the parameter values that merit a more careful study. The most reliable way to explore these complex dynamics comprehensively is through direct $N$-body simulations. Thus, the favoured parameters identified here serve as robust initial conditions for future, large-scale numerical investigations aimed at uncovering the precise role of IMBHs in shaping the distribution of stars in the Galactic centre.
@@ 6 Summary and Discussion
We have pointed out a novel phenomenon in multi-body gravitating systems subject to a predominantly spherical potential in which the bodies execute planar orbits. In these systems the orbital planes, represented by the angular momentum vectors, reorient rapidly due to vector resonant relaxation (VRR). We have shown that in these systems angular momentum vectors may exhibit bound long-lived sub-clusters in angular momentum space, which evolve together due to their strong VRR coupling. In particular, we examined two bodies on nearly coplanar orbits with initially nearly parallel angular momentum vectors, that initially precess about their net angular momentum, and experience perturbations from a distant fourth body whose angular momentum dominates the system. We derived the conditions under which the initially coupled pair becomes disrupted by the perturber, causing their angular-momentum vectors to precess independently around the perturber’s angular-momentum vector. We investigated various orbital configurations, including circular orbits (Sec. 3.1), eccentric radially non-overlapping orbits, and eccentric radially overlapping orbits (Sec. 3.3). A general stability criterion (15) was established and summarised by Eq. (19), validated through numerical simulations, as shown in Fig. 2. Additionally, we derived analytical expressions in the quadrupole approximation and for cases involving radially overlapping orbits. These analytical results were further numerically tested using the N-Ring code, simulating the double-orbit-averaged equations of motion in the VRR formalism, and using the direct $N$-body code $\phi$-CPU (Fig. 5). Comparing the results obtained with these different methodologies, we quantified the validity of the VRR formalism when applied to more realistic settings (Figs 6 and 7).
We applied this stability criterion to construct a simplified model representing the disc of young massive stars in the Galactic centre interacting with a massive external perturber, such as a putative IMBH or the circumnuclear gaseous torus (the circumnuclear disc). Our model provides order-of-magnitude estimates for a possible perturber’s orbital parameters required to produce specific observational features in the angular-momentum distribution of disc stars, notably the clustering of angular momentum vectors and forming a retrograde structure. These results also identify the possible parameters that can serve as initial conditions for large-scale direct $N$-body simulations of the Galactic centre, analogous to those performed by Panamarev & Kocsis (2022).
@@ 7 Acknowledgements
We are grateful to Peter Berczik for providing us with the CPU version of $\phi$-GPU. This work was supported by the Science and Technology Facilities Council Grant Number ST/W000903/1, and by a Leverhulme Trust International Professorship Grant (No. LIP-2020-014). We acknowledge the support of the Science Committee of the Ministry of Education and Science of the Republic of Kazakhstan (Grant No. AP22787256). The work of Y.B.G. was partly supported by a Simons Investigator Award to A.A. Schekochihin.
@@ Appendix A Legendre sums
The VRR Hamiltonian in Eq. (2) leads to equations of motion (Eq. (7)) containing infinite sums of the form $\sum_{\ell}P^{2}_{\ell}(0)P_{\ell}(x)\,z^{\ell}$, where $P^{2}_{\ell}(0)$ is given by Eq. (6). Using Stirling’s approximation to the Gamma function, one can show that for large $\ell$
@@
We will be dealing with the sums of the form $\sum_{\ell}\ell^{-n}P_{\ell}(x)\,z^{\ell}$. For $|z|\leq 1$ (i.e. for both non-overlapping and overlapping orbits) these sums admit closed-form expressions via the standard Legendre generating function. We briefly outline the method here:
@@ Appendix B Asymptotics of Vector Resonant Relaxation
Kocsis & Tremaine (2015) has shown that asymptotically for $\ell\rightarrow\infty$ (see Eq. B68 therein)
$\mathcal{J}_{ij\ell}=\frac{\mathcal{J}_{ij}}{\ell^{2}}~{}~{}~{}{\rm where}~{}~{}~{}\mathcal{J}_{ij}=\left\{\begin{array}[]{ll}\dfrac{Gm_{i}m_{j}}{\pi^{2}}\dfrac{[(1+e_{\rm in})(1-e_{\rm out})]^{3/2}}{(e_{i}e_{j})^{1/2}}\dfrac{r_{\rm a,in}^{\ell}}{r_{\rm p,out}^{\ell+1}}&{\rm if~{}non-overlapping}\,,\\[8.61108pt] \dfrac{4}{\pi^{3}}\dfrac{Gm_{i}m_{j}}{a_{i}a_{j}}I_{2}&{\rm if~{}overlapping}\,,\end{array}\right.$ (57)
@@
where $x=\cos\theta=\bm{L}_{i}\cdot\bm{L}_{j}$, $s=\sqrt{(1-x)/2}=\sin(\theta/2)$, $c=\sqrt{(1+x)/2}=\cos(\theta/2)$, $b_{1}=-\pi^{2}/24$, $b_{2}=2\chi(2^{-1/2})+\frac{1}{2}(\ln 2)^{2}-\frac{1}{8}\pi^{2}$, and $\chi(z)=\sum_{n=0}^{\infty}z^{2n+1}/(2n+1)^{2}$ is the Legendre-chi function which may be expressed with the dilogarithm $\mathrm{Li}_{2}(z)=\sum_{n=1}^{\infty}z^{n}/n^{2}$ as $\chi(z)=\frac{1}{2}[\mathrm{Li}_{2}(z)-\mathrm{Li}_{2}(-z)]$. The last approximate equality is accurate to better than $0.5\%$ for all $\theta$. Note that $b_{1}$ can be set to 0 in practice, that term does not affect the evolution. Equation (58) shows that for overlapping orbits, the VRR interaction energy is minimised for parallel or anti-parallel orbits (corresponding to stable equilibria) and it is maximised for perpendicular orbits (unstable equilibria) (Takács & Kocsis, 2018).

### v2

@@ 5 Applications to the Galactic Centre
For all pairs, we set eccentricities $e_{s}=0.3\pm 0.03$, and we explore a range of semi-major axes $0.04\leq a\leq 0.5~\text{pc}$. This parameter set mimics that of the young stars in the central parsec of the Galaxy (von Fellenberg et al., 2022). For the perturber, we set the eccentricity to $e_{\bullet}=0.3$ but vary its semi-major axis. The discrete values chosen are: $a_{\bullet}=0.15$ pc, the projected distance of the IRS13 association, potentially hosting an IMBH (Tsuboi et al., 2017); $a_{\bullet}=0.5$ pc, the possible 3D distance of IRS13 (Tsuboi et al., 2020); and $a_{\bullet}=1.5$ pc, an estimated distance to the circumnuclear disc (e.g. Oka et al. 2011). Additionally, we vary the mass of the perturber from $2\times 10^{3}$ to $10^{5}~\,{\rm M}_{\odot}$, placing lighter perturbers closer to the SMBH to operate within the allowed range of masses for the putative IMBHs (Lommen and Backer, 2001; Gualandris and Merritt, 2009; Gualandris et al., 2010; Naoz et al., 2020; GRAVITY Collaboration et al., 2020; Gravity Collaboration et al., 2023).
An important parameter is the mass of the fragment (i.e. a single clump) relative to the IMBH mass; we thus choose two different clump masses: $m_{i}+m_{j}=100\,{\rm M}_{\odot}$ and $m_{i}+m_{j}=200\,{\rm M}_{\odot}$, with $m_{i}=m_{j}$ in both cases. The clump mass is set by the number of particles with the strongest VRR coupling coefficients; in general, it may be defined by the number of objects with overlapping and marginally overlapping orbits and may depend on the position of objects within the disc. However, in our toy model we adopt the two values above for simplicity. Finally, we define the innermost pairs as the ones with semi-major axes $0.04\leq a\leq 0.05$ pc. Having set all the parameters we can use the disruption criterion in Eq. (19) which tells us that the pairs with $m_{i}+m_{j}=100\,{\rm M}_{\odot}$ will be decoupled only for the IMBH of $5\times 10^{3}\,{\rm M}_{\odot}$ and for some outer perturbers with $a_{\bullet}=0.5$ and 1.5 pc. On the other hand, while clumps with $m_{i}+m_{j}=200\,{\rm M}_{\odot}$ are expected to remain coupled for all inner perturbers with the parameters listed above, they can be decoupled by the IMBH of $10^{4}\,{\rm M}_{\odot}$ with $a_{\bullet}=0.5$ pc. Additionally, if we were to include interactions between different fragments, any clump whose orbit overlaps or marginally overlaps with that of the IMBH would be decoupled from the innermost clumps according to the general condition in Eq. (15) (which must be used for large orbital separations between clumps). Because we ignore such fragment–fragment interactions in our toy model, we instead treat the innermost region as precessing independently about the IMBH and measure the inclination of all other clumps with respect to this innermost disc. Thus, it transpires that for this toy model, the stellar disc is expected to be stable in most of the parameter space – the clumps should be coupled to each other but at the same time they should be de-coupled with respect to the inner clumps. Therefore, by running N-Ring simulations we can measure relative inclinations of the clumps with respect to the inner disc and relative inclination between the clumps at 10 Myr.
We ran a set of isolated three-ring problems in N-Ring with initial conditions adjusted to mimic the Galactic centre. We initialised all pairs with relative inclinations of $\theta_{b}=10^{\circ}$ and inclination with respect to the perturber $\theta_{\bullet}\simeq 65^{\circ}$. We chose this inclination because, in the case of disruption, the maximum relative angle $\theta_{1}+\theta_{2}=130^{\circ}$ exceeds $100^{\circ}$, pushing the corresponding pairs into retrograde orbits relative to the inner objects. For each set of initial conditions, we perform $10^{4}$ three-ring simulations.
@@
To interpret the results, we stacked all pairs in a single snapshot at $t=10~\text{Myr}$ and measured their relative inclinations $\theta_{b}$, as well as the relative inclination $\theta_{\mathrm{rel}}$ of each pair’s total angular momentum $\bm{L}_{i}+\bm{L}_{j}$ with respect to the net angular momentum of the innermost pairs (semi-major axes $0.04\leq a\leq 0.05$ pc).
@@
The simulation results indicate that, for clump masses of $100\,{\rm M}_{\odot}$, an IMBH of mass $5\times 10^{3}\,{\rm M}_{\odot}$ at $a_{\bullet}=0.15~\text{pc}$ pushes nearby pairs into retrograde orbits relative to the inner disc (orange curve, upper left panel) and significantly disrupts this region (orange curve, lower left panel). However, a stellar disc with clump masses of $200\,{\rm M}_{\odot}$ maintains its thickness (lower right panel, nearly constant lines), while simultaneously re-orienting regions around $0.15~\text{pc}$ into retrograde orbits. A lighter IMBH at this distance can reorient the disc less dramatically, reaching about $50^{\circ}$ relative inclination within 10 Myr. These cases resemble the observed distributions of young stars, notably in Bartko et al. (2009) – the low mass clumps produce many particles outside of the disc plane while the more massive clumps tend to preserve the original thickness of the disc. This suggests that in more realistic scenarios characterized by a variable stellar mass function and diverse orbital configurations, it is natural to expect a superposition of these effects resulting in complex disc structures, as recently demonstrated in observational studies (von Fellenberg et al., 2022; Jia et al., 2023).
@@
In summary, our toy model favours an IMBH mass of about $5\times 10^{3}\,{\rm M}_{\odot}$. However, this model treats the disc fragment as two rings, neglecting its interaction with the rest of the disc, and also does not include torques due to fluctuating anisotropy in the nuclear star cluster, which could be significant on the relevant time-scales (e.g. Panamarev and Kocsis 2022), nor does it account for the inward migration of the IMBH induced by Chandrasekhar dynamical friction (Chandrasekhar, 1943), or an orbital alignment due to resonant dynamical friction (Szölgyén et al., 2021; Ginat et al., 2023). Let us estimate the effect that dynamical friction with the nuclear cluster would have on the IMBH, and see if it could cause it to lose energy and sink into the SMBH. The (non-resonant) Chandrasekhar dynamical friction force on an IMBH of mass $m_{\bullet}$ moving with velocity $\mathbf{v}$ through a background of stars of density $\rho$ and one-dimensional velocity dispersion $\sigma$ can be written as
$\mathbf{F}_{\rm df}=-\,4\pi\,G^{2}\,m^{2}_{\bullet}\,\ln\Lambda\,\frac{\rho}{v^{3}}\,f(X)\,\mathbf{v}\,$, (39)
where $v=|\mathbf{v}|$, $\ln\Lambda$ is the Coulomb logarithm, $X\equiv v/(\sqrt{2}\,\sigma)$, and $f(X)$ is a dimensionless function of order unity that encodes the dependence on the ratio of the IMBH speed to the stellar velocity dispersion (Binney and Tremaine, 2008). For a circular Keplerian orbit of radius $a_{0}$ around the SMBH, with $v\simeq\sqrt{GM/a_{0}}$ and $\sigma\simeq v/\sqrt{2}$, this force implies a local dynamical-friction time-scale $t_{\rm df,loc}(a_{0})\equiv a_{0}/|\dot{a}|_{a_{0}}$ which can be compared to the local two-body relaxation time $t_{r}(a_{0})\simeq 0.34\,\sigma^{3}/(G^{2}m_{\star}\rho\ln\Lambda)$, where $m_{\star}$ is the typical stellar mass. Using these expressions ($f(X)\approx 0.428$) one finds
$t_{\rm df,loc}(a_{0})\simeq 0.77\frac{m_{\star}}{m_{\bullet}}t_{r}(a_{0}),$ (40)
so that for $m_{\bullet}\gg m_{\star}$ the IMBH inspirals on a time-scale much shorter than the local relaxation time. Taking $t_{r}\approx 10$ Gyr (see e.g. Kocsis and Tremaine 2011; Panamarev and Kocsis 2022) at the edge of the disc ($a_{0}\simeq 0.5$ pc), we find that for IMBHs with masses $m_{\bullet}\geq 10^{4}\,{\rm M}_{\odot}$ the dynamical friction time drops below $1$ Myr, which is significantly shorter than the age of the disc (661010 Myr; Bartko et al. 2010). We therefore conclude that such massive black holes are unlikely to exist within the disc.
Including the additional effects described above may reveal that lower-mass IMBHs can also effectively reorient the stellar disc into retrograde orbits within $\sim 10$ Myr. Although our simplified model provides useful order-of-magnitude estimates, it can neither rule out the existence of IMBHs, nor prove them – merely point to the parameter values that merit a more careful study. The most reliable way to explore these complex dynamics comprehensively is through direct $N$-body simulations. Thus, the favoured parameters identified here serve as robust initial conditions for future, large-scale numerical investigations aimed at uncovering the precise role of IMBHs in shaping the distribution of stars in the Galactic centre.
Additionally, gas forces may affect the dynamics at early times. The disc may become warped in the gas phase, before stars have formed (Levin and Beloborodov, 2003; Nealon et al., 2015), interaction of stars with the gaseous disc may lead to orbital alignment, circularization and semimajor axis decay (Kennedy et al., 2016; Panamarev et al., 2018) and/or the gas could affect the dynamics of the IMBH (Fabj et al., 2020). However, coupling gas physics with stellar dynamics requires computationally expensive hydrodynamical simulations integrated with $N$-body gravity, a task of significant complexity that demands a dedicated investigation and is therefore beyond the scope of this work.
@@ 6 Summary and Discussion
We have pointed out a novel phenomenon in multi-body gravitating systems subject to a predominantly spherical potential in which the bodies execute planar orbits. In these systems the orbital planes, represented by the angular momentum vectors, reorient rapidly due to vector resonant relaxation (VRR). We have shown that in these systems angular momentum vectors may exhibit bound long-lived sub-clusters in angular momentum space, which evolve together due to their strong VRR coupling. In particular, we examined two bodies on nearly coplanar orbits with initially nearly parallel angular momentum vectors, that initially precess about their net angular momentum, and experience perturbations from a distant fourth body whose angular momentum dominates the system. We derived the conditions under which the initially coupled pair becomes disrupted by the perturber, causing their angular-momentum vectors to precess independently around the perturber’s angular-momentum vector. We investigated various orbital configurations, including circular orbits (Sec. 3.1), eccentric radially non-overlapping orbits, and eccentric radially overlapping orbits (Sec. 3.3). A general stability criterion (15) was established and summarised by Eq. (19), validated through numerical simulations, as shown in Fig. 2. Additionally, we derived analytical expressions in the quadrupole approximation and for cases involving radially overlapping orbits. These analytical results were further numerically tested using the N-Ring code, simulating the double-orbit-averaged equations of motion in the VRR formalism, and using the direct $N$-body code $\phi$-CPU (Fig. 5). Comparing the results obtained with these different methodologies, we quantified the validity of the VRR formalism when applied to more realistic settings (Figs 7 and 8).
We applied this stability criterion to construct a simplified model representing the disc of young massive stars in the Galactic centre interacting with a massive external perturber, such as a putative IMBH or the circumnuclear gaseous torus (the circumnuclear disc). Our model provides order-of-magnitude estimates for a possible perturber’s orbital parameters required to produce specific observational features in the angular-momentum distribution of disc stars, notably the clustering of angular momentum vectors and forming a retrograde structure. These results also identify the possible parameters that can serve as initial conditions for large-scale direct $N$-body simulations of the Galactic centre, analogous to those performed by Panamarev and Kocsis (2022).
@@ 7 Acknowledgements
We thank the referee for useful comments and suggestions. We are grateful to Peter Berczik for providing us with the CPU version of $\phi$-GPU. This work was supported by the Science and Technology Facilities Council grant number ST/W000903/1, and by a Leverhulme Trust International Professorship Grant (No. LIP-2020-014). We acknowledge the support of the Science Committee of the Ministry of Education and Science of the Republic of Kazakhstan (Grant No. AP22787256). The work of Y.B.G. was partly supported by a Simons Investigator Award to A.A. Schekochihin.
@@ Appendix A Legendre sums
The VRR Hamiltonian in Eq. (2) leads to equations of motion (Eq. (7)) containing infinite sums of the form $\sum_{\ell}[P_{\ell}(0)]^{2}P_{\ell}(x)\,s_{ij\ell}\alpha^{\ell}$, where $P_{\ell}(0)]^{2}$ is given by Eq. (6). For circular orbits this is $\sum_{\ell}[P_{\ell}(0)]^{2}P_{\ell}(x)\,z^{\ell}$, where $z=r_{\mathrm{in}}/r_{\mathrm{out}}$. Kocsis and Tremaine (2015) showed (see their Appendix B5) that for both eccentric orbits
$s_{ij\ell}\alpha^{\ell}\rightarrow\frac{1}{\ell}\left(\frac{r_{\mathrm{a,in}}}{r_{\mathrm{p,out}}}\right)^{\ell}$ (41)
for large $\ell$. Using Stirling’s approximation to the Gamma function, one can show that for large $\ell$
@@
Thus, we will be dealing with the sums of the form $\sum_{\ell}\ell^{-n}P_{\ell}(x)\,z^{\ell}$. For $|z|\leq 1$ (i.e. for both non-overlapping and overlapping orbits) these sums admit closed-form expressions via the standard Legendre generating function. We briefly outline the method here:
@@
where $\chi(z)=\sum_{n=0}^{\infty}z^{2n+1}/(2n+1)^{2}$ is the Legendre-chi function which may be expressed with the dilogarithm $\mathrm{Li}_{2}(z)=\sum_{n=1}^{\infty}z^{n}/n^{2}$ as $\chi(z)=\frac{1}{2}[\mathrm{Li}_{2}(z)-\mathrm{Li}_{2}(-z)]$.
@@
A.1 Integral representations for fractional–power Legendre sums
The technique outlined above may be extended to evaluate sums containing fractional powers of $\ell$ in the denominator, such as
$\sum_{\ell}\frac{P_{\ell}(x)}{\ell^{3/2}}\,z^{\ell},\qquad\sum_{\ell}\frac{P^{\prime}_{\ell}(x)}{\ell^{3/2}}\,z^{\ell},$
which arise in the asymptotic treatment of the VRR Hamiltonian when using the Stirling approximation for $P_{2\ell}(0)]^{2}\!\propto\!\ell^{-1}$, yielding an overall $\ell^{-3/2}$ dependence in the coupling coefficients for the case of circular-eccentric orbital configurations (cf. Appendix B5 in Kocsis and Tremaine 2015). These series converge for all $|z|\leq 1$ and $|x|<1$ but do not possess closed–form elementary expressions; however, they admit rapidly convergent integral representations obtained from the Legendre generating function.
Starting from
$G(x,q)=\sum_{\ell=0}^{\infty}P_{\ell}(x)\,q^{\ell}=(1-2xq+q^{2})^{-1/2},\qquad|q|<1,$ (61)
we recall the general identity (valid for $\Re\,\nu>0$)
$\frac{1}{\ell^{\nu}}=\frac{1}{\Gamma(\nu)}\int_{0}^{1}(-\ln t)^{\nu-1}\,t^{\,\ell-1}\mathrm{d}t,$ (62)
which provides an integral kernel for replacing powers of $\ell$ by logarithmic weights. Applying this to the generating function gives the Abel transform representation
$\sum_{\ell=0}^{\infty}\frac{P_{\ell}(x)}{\ell^{\nu}}\,z^{\ell}=\frac{1}{\Gamma(\nu)}\int_{0}^{1}\frac{(-\ln t)^{\nu-1}}{t}\,G(x,zt)\,\mathrm{d}t,\qquad 0<z\leq 1.$ (63)
For $\nu=\tfrac{3}{2}$ this yields the expression used in our numerical evaluation:
$S_{1}(x,z)\equiv\sum_{\ell=0}^{\infty}\frac{P_{\ell}(x)}{\ell^{3/2}}\,z^{\ell}=\frac{2}{\sqrt{\pi}}\int_{0}^{1}\frac{\sqrt{-\ln t}}{t}\,G(x,zt)\,\mathrm{d}t.$ (64)
[…]
$G_{\mathrm{even}}(x,q)=\frac{1}{2}\left[G(x,q)+G(x,-q)\right]=\sum_{\ell\,\mathrm{even}}P_{\ell}(x)\,q^{\ell},$ (65)
therefore leads to
$S_{1}^{(\mathrm{even})}(x,z)=\frac{2}{\sqrt{\pi}}\int_{0}^{1}\frac{\sqrt{-\ln t}}{t}\bigl[G_{\mathrm{even}}(x,zt)-1\bigr]\,\mathrm{d}t=\frac{2}{\sqrt{\pi}}\int_{0}^{1}\frac{\sqrt{-\ln t}}{t}\left[\frac{1}{2}\!\left(\frac{1}{\sqrt{1-2xzt+z^{2}t^{2}}}+\frac{1}{\sqrt{1+2xzt+z^{2}t^{2}}}\right)-1\right]\mathrm{d}t.$ (66)
The subtraction of unity removes the $\ell=0$ term, ensuring convergence at the lower limit.
An entirely analogous procedure applies to the series containing the derivatives $P^{\prime}_{\ell}(x)$. Differentiating the generating function with respect to $x$ gives
$\frac{\partial G(x,q)}{\partial x}=\sum_{\ell=0}^{\infty}P^{\prime}_{\ell}(x)\,q^{\ell}=\frac{q}{(1-2xq+q^{2})^{3/2}}.$ (67)
[…]
$S_{2}^{(\mathrm{even})}(x,z)=\frac{1}{\sqrt{\pi}}\,z\int_{0}^{1}\!\sqrt{-\ln t}\,\Bigl[(1-2xzt+z^{2}t^{2})^{-3/2}-(1+2xzt+z^{2}t^{2})^{-3/2}\Bigr]\,\mathrm{d}t.$ (68)
Equations (66) and (68) provide rapidly convergent one–dimensional integrals that can be evaluated numerically with high precision for any $(x,z)$ in the physically relevant domain $0<x<1$, $0<z\leq 1$.
@@ Appendix B Asymptotics of Vector Resonant Relaxation
Kocsis and Tremaine (2015) has shown that asymptotically for $\ell\rightarrow\infty$ (see Eq. B68 therein)
[…]
@@
where $x=\cos\theta=\bm{L}_{i}\cdot\bm{L}_{j}$, $s=\sqrt{(1-x)/2}=\sin(\theta/2)$, $c=\sqrt{(1+x)/2}=\cos(\theta/2)$, $b_{1}=-\pi^{2}/24$, $b_{2}=2\chi(2^{-1/2})+\frac{1}{2}(\ln 2)^{2}-\frac{1}{8}\pi^{2}$, and $\chi(z)=\sum_{n=0}^{\infty}z^{2n+1}/(2n+1)^{2}$ is the Legendre-chi function which may be expressed with the dilogarithm $\mathrm{Li}_{2}(z)=\sum_{n=1}^{\infty}z^{n}/n^{2}$ as $\chi(z)=\frac{1}{2}[\mathrm{Li}_{2}(z)-\mathrm{Li}_{2}(-z)]$. The last approximate equality is accurate to better than $0.5\%$ for all $\theta$. Note that $b_{1}$ can be set to 0 in practice, that term does not affect the evolution. Equation (71) shows that for overlapping orbits, the VRR interaction energy is minimised for parallel or anti-parallel orbits (corresponding to stable equilibria) and it is maximised for perpendicular orbits (unstable equilibria) (Takács and Kocsis, 2018).